OpenStudy (moonlight):

solve the logarithmic log base6 (4x-9) = log base6 x

OpenStudy (moonlight):

please , thats my final exam ,,, need help

OpenStudy (anonymous):

I believe we can raise both sides to by 6 and thus end up with: 4x-9=x, next we subtract x from both sides giving us: 3x-9=0. then add 9 to both sides and finally divide both sides by three: x=3

OpenStudy (anonymous):

\[4x-9=x\] Next subtract x from both sides: \[3x-9=0\] Then: add 9 to both sides: \[3x=9\] Finally divide both sides by three: \[x=\]

OpenStudy (anonymous):

x=3*

OpenStudy (anonymous):

what do you think?

OpenStudy (moonlight):

thanx alot , i appreciated

OpenStudy (moonlight):

ur smart :)

OpenStudy (anonymous):

if you really wan to thank me, you should click the good answer button :)

OpenStudy (moonlight):

sure

OpenStudy (anonymous):

Thank you

OpenStudy (moonlight):

ur welcome, thank u thank u thank u